a. An airplane has a mass of 3.79 × 104 kg and takes off under the...

a.

An airplane has a mass of 3.79 × 104 kg and takes off under the influence of a constant net force of 3.45 × 104 N. What is the net force that acts on the plane's 82.7-kg pilot?

Number = ______ N

b.

A person in a kayak starts paddling, and it accelerates from 0 to 0.527 m/s in a distance of 0.452 m. If the combined mass of the person and the kayak is 64.4 kg, what is the magnitude of the net force acting on the kayak?

Number = _____ N
Answer #1

Part b)

We know, vf^2 =v0^2 + 2aΔx

Where, final velocity (vf) = .527 m/s

inital velocity (v0) = 0

Δx = .452 m

Now, lets find acceleration , a

.527^2=0^2+2*a*0.452

A=0.307 m/s^2

We also know,

F = ma

64.4 kg * .307 m/s^2 = 19.78 N (Ans)

=========